Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 14/3/2017

Location : CR21 in R1 at 2pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • DC update
  • List of Hall activities in the shutdown (Colin) : http://micewww.pp.rl.ac.uk/projects/operations/wiki/Hall_Task_List_Spring17
  • List of issues arising from previous user cycle (may overlap with Colin's list) :
    • Get a spare EMR board
    • Compressor 2 SSU communication problems (Ajit is aware)
    • Communication dropouts in the FCD helium level monitor
    • TKU compressor 2 needs servicing/ spares.
    • TKD may need to swap a waveguide to improve dead channel placement.
    • TKU butterfly helium leak tighting
    • Status of target1 backup machine in R78
    • need to be emphasise procedures to gain swipe access in the vac area for new shifters
    • need a spare PCI board for the tracker ioc pc backup machine
    • MLCR network isolation test still needs to be done
  • Controls and Monitoring
    • Lots of alarms in the previous run which were silenced "by default" e.g. Tracker cryo 2 / SSU
    • Are all the limits for each subsystem correctly defined?
  • Recurring list (may overlap with Colin's list)
    • #1878, #1891 Trench water pump control - pressure appears stable; shifter lockout in progress
    • #1881 Water system PVs not being archived - Services/Environment/Computing spreadsheet
    • #1880 Alarm Handler not reacting to conventional magnet interlock
    • #1903, #1881 micestore disk full, also needs to be in Nagios warning checks in place but need testing
    • #1881 File Compactor failure - shifters reported only the "previous error" message; should be in Alarm Handler. Move File Compactor output to top of RC screen? Also waiting on Services/Environment/Computing spreadsheet_
    • VME crates in Hall overheating; crate monitoring looking at wrong crate Services/Environment/Computing spreadsheet, GUI needs fixing
    • Is MOM DC Configuration Checklist Elog tab up-to-date?
    • webcams - images not visible on micewww, probably due to micenet lockout #1399 but low priority DONE
    • Neutron monitor & restart docs - MICE to take the CMS ION to Poole to be fixed ourselves. MLCR needs to know how to restart remote viewing of the temporary monitor.

AOB:

Minutes

  • CC Magnet status (CWhyte)
    Started the LH2 installation: installed absorber, applied MLI tape, finishing Indium seals this afternoon.
    The Solenoid-mode test for SSD has been done.
    The FC power supply went Phut TM while isolating. Tested damaged PSU with replaced fuses; still failed. Going for external repair - expect 7-14 days. Controller looks OK.
  • Hydrogen status (VBayliss)
    On schedule: Wed. (tomorrow) UPS test; Thu. H2 detector system calibration; next Mon.-Wed. PW coming down to work on controls.
    Presently installing absorber and turret; doing Indium seals and leak-checking.
    Have a contractor coming in to modify exhaust line.
  • Diffuser iris 4 (CWhyte)
    Are deciding if need one or both PRY end-plates out.
    Need to double-check TOF1 cabling before moving - it got re-connected wrongly last time.
    Get Joe to come out and look at it, and then decide: it may be possibly to access that iris (the most UPstream) without fully removing the Diffuser assembly. If the diffuser assembly comes out then also need to put out lights/put up tenting to protect Tracker...
  • C&M (AKurup)
    Is working through restarting IOCs
    Needs to restart Archiver (miceecserv) on Thursday and also ALH to get configuration update.
    Water pump system has been passworded - needs to send password to Henry
  • Detectors
    Have swapped to the other spare compressor (CSW-71 instead of F70) on Tracker 2, temperatures now OK. Still trying to get the Sumitomo rep in to get the heads serviced.

AOB:

  • We are to get the CMS-ION neutron detector back - need it before DSA shuts.
  • On the last night of running, Run Control managed to spawn two simultaneous File Compactor processes, which were then left running during the rest of the data-taking. Has anyone checked whether or not subsequent data was affected?
    Durga has checked the data [although he has only checked that it unpacks correctly, not for any other likely side-effects such as an increase in effective dead-time.]

Actions

  • JBoehm to get SS operating info from PHanlet (ongoing)
  • JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI"
  • JBoehm arrange tests of polarity effect on SSD leaks
  • Hall Manager when removing FC to make a crude check that LiH hasn't moved.
  • Hall Manager: water system change-over procedure
  • Status of miceraid5 disk cleanup (stands)
  • DC to ensure Health Physics have checked temporary neutron monitor (stands)
  • SBoyd - ensure list of issues arising from current running is being collated
  • CWhyte - draw up the shutdown task list
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel
  • PFranchini - Get all shifters to go through the CE checklist during Control Room refresher training
  • MUchida - investigate moving Tracker waveguides around to cure "hole"
  • AKurup - send water pump GUI password to Henry

Updated by Nebrensky, Henry over 6 years ago ยท 15 revisions